OpenStudy (anonymous):

A standard number cube is tossed. Find P(even or greater than 4).

OpenStudy (anonymous):

Desired outcomes: 2,4,5,6 All outcomes: 1,2,3,4,5,6 # of desired outcomes: 4 Total # of possible outcomes: 6\[P = \frac{\textrm{# of desired outcomes}}{\textrm{total # of outcomes}} = \frac{4}{6} = \frac{2}{3}\]

OpenStudy (anonymous):

thanks